For the following set of scores, find the value of each expression: Σ X Σ 3*X2...

  1. For the following set of scores, find the value of each expression:
    1. Σ X
    2. Σ 3*X2
    3. Σ (X-3)
    4. Σ (7X-10)2


X  
18
5
-9
11
9
-14
-3
Answer #1

\sum X = [ 18 + 5 - 9 + 11 + 9 - 14 - 3) = 17

\sum 3 * X2 = 3 * 182 + 3 * 52 + 3 * (-9)2 ​​​​​​​+3 * 112 ​​​​​​​+3 * 92 ​​​​​​​+3 * (-14)2 ​​​​​​​+3 * (-3)2 ​​​​​​​= 2511

\sum ( X - 3) = (18-3) + (5 -3) + (-9 -3) + (11 -3) + (9 -3) + (-14 -3) + (-3 -3) = -4

\sum ( 7X -10)2 = (7 *18 - 10)2 + (7 *5 - 10)2​​​​​​​ + (7 *(-9) - 10)2​​​​​​​ + (7 *11 - 10)2​​​​​​​ + (7 *9 - 10)2​​​​​​​ + (7 *(-14) - 10)2​​​​​​​ + (7 *(-3) - 10)2​​​​​​​

= 39333